OpenStudy (anonymous):

Which expression correctly shows the Commutative property? 8 + 14 + 22 + 9 = (8 + 14) + (22 + 9) 8 + 14 + 22 + 9 = 8 + (14 + 22) + 9 8 + 14 + 22 + 9 = 53 8 + 14 + 22 + 9 = 8 + 22 + 14 + 9

OpenStudy (freckles):

commutative just means the order has changed for example a+b=b+a is what the commutative property over addition looks like

OpenStudy (freckles):

what is the only choice where the order has changed

OpenStudy (anonymous):

the 1st one

OpenStudy (freckles):

lets think about it what does order mean exactly like what i mean by order is that the terms get switched around and not all the terms have to be switched around but at least some do in order for it be consider commutative property

OpenStudy (freckles):

example of commutative property being applied are: 8+7+5=7+8+5

OpenStudy (freckles):

a+b+c=b+a+c <--commutative property has been applied the order of the terms has changed

OpenStudy (freckles):

a+b+c=a+c+b see the second term has become the new third term and the third term has become the new 2nd term

OpenStudy (anonymous):

the 2nd one

OpenStudy (freckles):

order has not changed there

OpenStudy (freckles):

8 + 14 + 22 + 9 = 8 + 22 + 14 + 9 look at this one this is exactly what I have been saying the 14 and 22 have switched places the order has been changed here has it not?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so its the last one right

OpenStudy (freckles):

if you are convinced it is so, then I will agree with that

OpenStudy (anonymous):

because the 14 +22 is changed

OpenStudy (freckles):

yep yep

OpenStudy (freckles):

we had 14+22 became 22+14